OpenStudy (anonymous):

A sample of 100 janitors in a certain school district found that the average time each janitor had worked for the school district was 8 years. Suppose we know that the standard deviation of the population is s=5 years. Construct a 95% confidence interval for the mean time of the population of janitors have spent with the school district. 8±1.98 8±0.82 8±0.71 8±1.53 8±0.98

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@poopsiedoodle

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

@beccaboo333

OpenStudy (beccaboo333):

Sorry, I don't know :/

OpenStudy (anonymous):

?

OpenStudy (nikato):

I don't really know, but you can see if this helps http://m.wikihow.com/Calculate-Confidence-Interval

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

A.?

OpenStudy (anonymous):

help guys!

OpenStudy (anonymous):

Thank you @ybarrap

OpenStudy (ybarrap):

you're welcome

OpenStudy (anonymous):

D.?

OpenStudy (anonymous):

okay last one

OpenStudy (anonymous):

The same sample of 25 seniors from the urban school district with the mean and standard deviation N(450, 100). A 95% confidence interval for µ for the population of seniors with a margin of error of ± 25 is used. What is the smallest sample size we can take to achieve this same margin of error?

OpenStudy (anonymous):

@agent0smith your turn! :)

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@phi

OpenStudy (ybarrap):

I think this might be more accurate since the samples are so high: http://www.wolframalpha.com/input/?i=confidence+interval+mean&a=FSelect_**ZInterval-.dflt-&f2=0.95&f=ZInterval.c%5Cu005f0.95&f3=100&f=ZInterval.n%5Cu005f100&f4=5&f=ZInterval.sigma_5&f5=8&f=ZInterval.xbar%5Cu005f8 This uses the z-score rather than the student's T distribution with are based on small samples.

OpenStudy (ybarrap):

This means $$ \bar{x}=8\pm0.98 $$

OpenStudy (anonymous):

well here are my choices and i say it would be 62? 26 45 50 62 69

OpenStudy (ybarrap):

(I was still on the first, lol)

OpenStudy (anonymous):

OHh Sorry! yes I knew about that! isaw the link

OpenStudy (anonymous):

thats why i said thanks before lol

OpenStudy (anonymous):

I know from the first question that my confidence interval would be 8 plinus 0.98

OpenStudy (anonymous):

but i need help on this other quesiton The same sample of 25 seniors from the urban school district with the mean and standard deviation N(450, 100). A 95% confidence interval for µ for the population of seniors with a margin of error of ± 25 is used. What is the smallest sample size we can take to achieve this same margin of error? 26 45 50 62 69

OpenStudy (ranga):

Margin of error is inversely related to the square root of the sample size.

OpenStudy (anonymous):

@ranga i already no that, you're no help

OpenStudy (anonymous):

*know

OpenStudy (ybarrap):

(Please be polite when someone is trying to guide you through these problems) Confidence interval is, $$ \large \mu\pm\cfrac{z_{(1-c)\sigma}}{\sqrt n} $$ We want margin of error to be \(\pm25\), so $$ \large{ 25=\cfrac{z_{(1-c)/2}\sigma}{\sqrt n}\\ \sqrt n=\cfrac{z_{(1-c)/2}\sigma}{25}\\ n=\left (\cfrac{z_{(1-c)/2}\sigma}{25}\right )\\ } $$ Where \(c=0.95\) and \(\sigma=100\). This is the smallest sample size that will satisfy the requirement. We now just need \(\large z_{(1-c)/2}=z_{0.025}\) Do you know how to find this?

OpenStudy (anonymous):

@ybarrap ?

OpenStudy (anonymous):

hmm yes hold on

OpenStudy (anonymous):

50

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

We need \(\large z_{0.025}\) to plug in the equation above, I'd have to look it up, I thought you were doing that.

OpenStudy (anonymous):

i did......

OpenStudy (ybarrap):

what was it?

OpenStudy (anonymous):

my calculator does it for me so i dont know, i have a n-spire im just saying overall it is c

OpenStudy (anonymous):

because i plugged in those numbers into that equation you gave me

OpenStudy (ybarrap):

ok, let me look it up

OpenStudy (anonymous):

alright

OpenStudy (ybarrap):

I need to go, I'll check back later if you still need validation

OpenStudy (anonymous):

@agent0smith :(

OpenStudy (anonymous):

I've been on this question for the past hour please help me

OpenStudy (anonymous):

i was saying that it was 50?